A toy is hanging at rest at the end of a spring. The toy is 350 g, and stretches the spring

10 cm. What is the spring constant?


1
Expert's answer
2021-05-20T10:08:40-0400
"F=kx,""mg=kx,""k=\dfrac{mg}{x}=\dfrac{0.35\ kg\cdot9.8\ \dfrac{m}{s^2}}{0.1\ m}=34.3\ \dfrac{N}{m}."